On September 30, 2021, Bricker Enterprises purchased a machine for $220,000. The estimated service life is 10 years with a $26,000 residual value. Bricker records partial-year depreciation based on the number of months in service.

Depreciation for 2021, using the double-declining-balance method, would be: (Do not round intermediate calculations.)

Multiple Choice

  • $44,000.

  • $11,000.

  • $38,800.

  • $9,700.
Answer #1

Correct answer-----------$11,000

Working

Double declining Method
A Cost $ 220,000
B Residual Value $ 26,000
C=A - B Depreciable base $ 194,000
D Life [in years] 10
E=C/D Annual SLM depreciation $ 19,400
F=E/C SLM Rate 10.00%
G=F x 2 DDB Rate 20.00%

.

Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
2021 $ 220,000 20.00% $ 11,000* $ 11,000 $ 209,000

*For three months {(220000 x 20% ) x 3/12}